# Remote Sensing Spatio-Temporal Vision-Language Models: A Comprehensive Survey

### v2

@@ II-B1 Remote Sensing Change Captioning
Current spatio-temporal vision-language understanding research primarily focuses on the remote sensing change captioning (RS-CC) task. This task aims to generate detailed, natural language descriptions that explain the differences between multi-temporal remote sensing images. This process not only depends on precise visual change recognition but also demands high language generation capability to ensure both accuracy and fluency of language [25, 26, 91, 92, 93, 94, 95, 96, 97, 98]. RS-CC methods typically adopt a three-stage architecture as illustrated in Fig. 4 (a).
In the visual encoding stage, bi-temporal or multi-temporal images are processed by Siamese encoders—often implemented using Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs)—to extract rich spatial features from each image. The next stage, spatio-temporal feature integration, is critical for integrating the extracted features from different time points. In this stage, the model employs various techniques, such as multi-scale feature fusion strategies and attention mechanisms, to identify spatio-temporal change patterns and highlight significant changes while suppressing irrelevant pseudo-changes caused by lighting or weather conditions [71, 73, 27]. For instance, Liu et al.[26] introduced a Transformer-based method with cross-encoding modules that effectively leverage differential features to focus on changing regions, while Chang et al.[25] proposed a hierarchical self-attention network to dynamically emphasize spatio-temporal changes.
@@
Pix4Cap [100] 2024.01 ViT-B/32 Tr-Dec Utilized a BIT model pre-trained on the LEVIR-CD dataset to establish CD pseudo-labels for the LEVIR-CC dataset, introducing a change detection branch in the change caption network to learn pseudo-labels.
Change-Agent [101] 2024.03 ViT-B/32 Tr-Dec Built a change agent comprising an LLM and a multi-level change interpretation model. Proposed a dual-branch multi-task model and LEVIR-MCI dataset containing masks and descriptions. [Link]
Semantic-CC [102] 2024.07 SAM Vicuna Semantic-CC leveraged the latent knowledge of the SAM foundational model to alleviate the high generalization algorithm’s dependence on extensive annotations and generated more comprehensive and accurate change descriptions through pixel-level semantic guidance from change detection.
DetACC [103] 2024.09 ResNet-101 Tr-Dec Enhanced its change description capabilities by utilizing explicit visual change information.
KCFI [104] 2024.09 ViT Qwen Proposes a novel multimodal framework for remote sensing change captioning, guided by key change features and instruction tuning (KCFI). Extracted key change features and leveraged LLM to improve the accuracy of the description. Using dynamic weights to optimize the losses for CC and CD. [Link]
ChangeMinds [105] 2024.10 Swin Transformer Tr-Dec Proposed ChangeLSTM module that utilizes the recent XlSTM to process features from both directions to obtain a universal change-aware representation. Proposed the Multi-task Predictor that uses the unified change decoder to fuse change-aware representations and employs multi-task classifiers to generate predictions for both CD and CC tasks. [Link]
CTMTNet [106] 2024.10 ResNet-101 Tr-Dec Proposed a dual-branch architecture, designed a Multi-Attention Feature Enhancement Module and Feature Fusion Block to enhance location perception.
MModalCC [93] 2025.01 ResNet-101 Tr-Dec Introduced SECOND-CC dataset featuring high-resolution RGB image pairs, semantic segmentation maps, and real-world scenarios, which contains 6 041 pairs of bi-temporal RS images and 30 205 sentences describing the differences. And proposed MModalCC, a multimodal framework that integrates semantic and visual data using advanced attention mechanisms. [Link]
Mask Approx Net [95] 2025.02 ResNet Tr-Dec Proposed a diffusion model-based approach Mask Approx Net, transitioning from feature learning to data distribution learning. It includes a multi-scale change detection module and a frequency-guided complex filter module. [Link]
FST-Net [107] 2025.04 Segformer-B1 Tr-Dec Proposed the FST-Net, using Frequency–Spatial Fusion module implements adaptive spectral decomposition to disentangle structural changes from high-frequency-noise artifacts. a unified dual-task decoder was used to achieve joint optimization of pixel-level change detection and semantic-level change captioning.
@@ II-B2 Multi-task learning of Change Detection and Change Captioning
Change detection and change captioning address complementary aspects of remote sensing spatio-temporal vision-language understanding. Change detection focuses on producing pixel-level change masks to pinpoint changed areas, whereas change captioning aims to generate semantic-level descriptions that reveal object attributes and contextual relationships. Recognizing the intrinsic connection between these tasks, recent research has integrated them into a unified multi-task learning framework to enhance overall efficiency and accuracy [100, 102, 104, 105, 106]. Some representative methods are summarized in Table II.
One notable approach, Change-Agent [101], laid the groundwork by introducing a typical framework as shown in Fig. 4 (b). It employs a shared visual encoder to extract spatio-temporal features from remote sensing temporal images, followed by two task-specific branches: one for change detection and another for change captioning. In this framework, the change detection branch leverages multi-scale features to generate precise and detailed change masks, while the captioning branch typically utilizes deeper visual features to capture the semantics of the object changes. Such a design closely resembles single-task models in each branch yet benefits from joint optimization.
@@
change-aware VQA [110] 2022.07 CNN RNN First explored this task and proposed a dataset made of multi-temporal image-question-answer triplets and a baseline method for this task.
CDVQA-Net [28] 2022.09 CNN RNN Build a dataset and a baseline containing multi-temporal feature encoding, multi-temporal fusion, multi-modal fusion, and answer prediction. [Link]
ChangeChat [111] 2024.09 CLIP-ViT Vicuna-v1.5 The first bi-temporal VLM specifically designed for remote sensing change analysis; ChangeChat leverages multimodal instruction tuning to handle complex queries; Developed a ChangeChat-87k dataset. [Link]
CDChat [112] 2024.09 CLIP ViT-L/14 Vicuna-v1.5 Constructed a 19K dataset and finetuned the LLaMA-1.5 model on the change captioning instruction dataset. [Link]
TEOChat [68] 2024.10 CLIP ViT-L/14 LLaMA-2 Proposed a VLM TEOChat for temporal EO data and a 554K temporal EO dataset. Experimented on multiple temporal tasks, outperforming video-Llava. [Link]
GeoLLaVA [113] 2024.10 Video encoder LLaVA-NeXT, Video-LLaVA Introduced a novel dataset and applied fine-tuning techniques to enhance VLM for detecting temporal changes in geographical landscapes. Enhanced models like Video-LLaVA and LLaVA-NeXT-Video use methods such as LoRA and QLoRA. [Link]
VisTA [29] 2024.10 Shared CLIP Image Encoder CLIP Text Encoder Introduced a new task called Change Detection Question Answering and Grounding (CDQAG). Constructed the first CDQAG benchmark dataset, QAG-360K, containing over 360K triplets of questions, text answers, and corresponding high-quality visual masks. Proposed a SOTA baseline on CDQAG dataset. [Link]
RSUniVLM [114] 2024.12 Siglip-400m Qwen2-0.5B Proposed RSUniVLM, a unified end-to-end remote sensing vision-language model for comprehensive multi-granularity understanding. It features a novel Granularity-oriented Mixture of Experts architecture, constrained to 1B parameters, and proposed a large-scale instruction dataset. [Link]
EarthDial [115] 2024.12 InternViT-300M Phi-3-mini Proposd EarthDial, a conversational VLM for multi-spectral, multi-temporal, and multi-resolution EO imagery. Introduced a large remote sensing instruction tuning dataset with 11.11M pairs.
UniRS [116] 2024.12 Siglip-400m Sheared-LLAMA-3B proposd UniRS, a VLM for multi-temporal remote sensing tasks, establishing a unified framework for single images, dual-time pairs, and videos. Introduced a Change Extraction module for dual-time images, a prompt augmentation mechanism for inference, and a multi-task joint fine-tuning framework. [Link]
@@ II-B3 Remote Sensing Change Question Answering
In the visual encoding stage, the visual encoder is often employed to extract features from multi-temporal images, with the resulting features fused to capture dynamic visual information. In the question encoding stage, pre-trained language models (e.g., BERT [117] or GPT [118]) transform the user’s query into a semantic embedding suitable for model understanding. During multimodal interaction, attention mechanisms (e.g., self-attention and cross-attention) are widely applied to align visual features with the linguistic cues provided by the query, allowing the model to focus on the most relevant spatial regions. Finally, in the answer generation stage, the fused multimodal representation is converted into a natural language response.
@@ II-B4 Remote Sensing Text-driven Temporal Images Retrieval
One of the primary challenges in RSI-TIR is the issue of false negatives. Specifically, an image pair labeled as a negative sample in a training batch may actually be a positive sample that matches the query text, thus hindering effective model training. This problem is common in many contrastive learning tasks and can be addressed by some solutions [124, 125]. To mitigate the issue in RSI-TIR, Ferrod et al. [79] implemented two strategies: False Negative Elimination (FNE), which excludes potential false negatives from the loss computation to prevent interference, and False Negative Attraction (FNA), which re-labels these samples as positive to better align with the true relationships in the data. These techniques improve retrieval precision in complex spatio-temporal scenarios.
@@ II-B5 Remote Sensing Change Grounding
Methods such as TEOChat [68] and EarthDial [115] adopt bounding box outputs. For instance, TEOChat [68] use a model architecture inspired by LLaVA-1.5 [119], which utilizes a temporally shared ViT-L/14 to encode the temporal images. The resulting embeddings are mapped via an MLP and fed into LLaMA-2 [126], with the language model generating bounding box coordinates in textual format—thus effectively grounding the detected changes according to the input query.
On the other hand, methods that produce mask outputs include VisTA [29], ChangeChat [111], RSUniVLM [114], Falcon [127], and GeoRSMLLM [128]. For example, Li et al. [29] introduced VisTA, a multi-task model designed for change question answering and grounding. VisTA simultaneously generates a pixel-level change mask and a textual answer. The textual response is produced via a two-layer MLP, while the mask decoder employs two attention blocks, enabling the model to provide both semantic and visual explanations. This dual-output approach exemplifies a versatile solution for RS-CG tasks, enhancing the overall interpretability and precision of change localization in remote sensing imagery.
@@
III Key Technologies in Spatio-Temporal Visual-Language Models
III-A Preliminary Knowledge of Basic Models
III-A1 LSTM
The Long Short-Term Memory (LSTM) [129] is a specialized type of Recurrent Neural Network (RNN) designed to capture long-term dependencies in sequential data. Traditional RNNs suffer from vanishing or exploding gradients during long sequence modeling, a problem that LSTM addresses by introducing memory cells and gating mechanisms.
@@ III-A1 LSTM
where $x_{t}$ is the current input, $h_{t-1}$ is the previous hidden state, and $\sigma$ is the sigmoid function.
@@
This architecture enables LSTM to capture long-term dependencies, making it widely applicable in fields such as natural language processing, speech recognition, and time series forecasting. Recent extensions like xLSTM [130] further refine these capabilities for more complex tasks.
Figure 6: The architecture of Transformer [131].
III-A2 Transformer
@@
Currently, the Transformer has been adapted in many fields, including NLP, computer vision [132, 133], multimodality [134, 135], medicine [136, 137], and remote sensing [138, 139]. Transformers’ advanced scalability has promoted the emergence of various state-of-the-art LLMs, such as chatgpt [140], T5 [141], LLaMA2 [126], and Qwen [142].
Figure 7: The architecture of Mamba [99].
III-A3 Mamba
@@
Mamba extends original SSMs by introducing a selective state space mechanism, allowing parameters like $\bar{\mathbf{B}}$, ${\mathbf{C}}$, and $\mathbf{\Delta}$ to adapt dynamically to the input sequence. This adaptability enables the model to selectively attend to different parts of the input, making it contextually aware of processing sequences in an efficient manner. Due to the global receptive field and linear computational complexity, Mamba [99] has demonstrated impressive performance across various tasks [144, 145, 146, 147, 148].
Figure 8: The key technologies can be applied across different remote sensing spatio-temporal vision-language tasks, including temporal image encoding, spatio-temporal visual feature fusion, language encoding, vision-language interaction, and language generation.

### v3

@@ II-B1 Remote Sensing Change Captioning
Current spatiotemporal vision-language understanding research primarily focuses on the remote sensing change captioning (RS-CC) task. This task aims to generate detailed, natural language descriptions that explain the differences between multi-temporal remote sensing images. This process not only depends on precise visual change recognition but also demands high language generation capability to ensure both accuracy and fluency of language [27, 28, 117, 118, 119, 120, 121, 122, 123, 124]. Previous RS-CC methods typically adopt a three-stage architecture as illustrated in Fig. 4 (a).
In the visual encoding stage, bi-temporal or multi-temporal images are processed by Siamese encoders, often implemented using Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs), to extract rich spatial features from each image. The next stage, spatiotemporal feature integration, is critical for integrating the extracted features from different time points. In this stage, the model employs various techniques, such as multi-scale feature fusion strategies and attention mechanisms, to identify spatiotemporal change patterns and highlight significant changes while suppressing irrelevant pseudo-changes caused by lighting or weather conditions [97, 99, 29]. For instance, Liu et al.[28] introduced a Transformer-based method with cross-encoding modules that effectively leverage differential features to focus on changing regions, while Chang et al.[27] proposed a hierarchical self-attention network to dynamically emphasize spatiotemporal changes.
@@
Pix4Cap [126] 2024.01 ViT-B/32 Tr-Dec Utilized a BIT model pre-trained on the LEVIR-CD dataset to generate pseudo-labels for the LEVIR-CC dataset. Introduced a change detection branch to guide the captioning network via pseudo-label supervision.
Change-Agent [127] 2024.03 ViT-B/32 Tr-Dec Built LEVIR-MCI dataset containing masks and captions, proposed a dual-branch multi-task framework for change detection and captioning, and designed a Change-Agent composed of an LLM and a multi-level interpretation model. [Link]
Semantic-CC [128] 2024.07 SAM Vicuna Leveraged the latent knowledge of the SAM foundation model and employed pixel-level semantic guidance from change detection to generate more comprehensive and accurate change descriptions.
DetACC [129] 2024.09 ResNet-101 Tr-Dec Enhanced changed regions by explicitly incorporating visual change detection masks to guide caption generation.
KCFI [130] 2024.09 ViT Qwen Following the dual-branch multi-task framework, it utilizes LLMs to generate more accurate descriptions and introduced dynamic weights strategy to balance detection and captioning losses. [Link]
ChangeMinds [131] 2024.10 Swin Transformer Tr-Dec Proposed an XlSTM-based ChangeLSTM module to process bi-temporal features from two directions to obtain a universal change-aware representation, followed by the multi-task predictors to predict change masks and captions. [Link]
CTMTNet [132] 2024.10 ResNet-101 Tr-Dec Constructed a WHU-CDC dataset containing binary building change masks and text descriptions and designed a Multi-Attention Feature Enhancement Module and Feature Fusion Block to enhance change location perception.
MModalCC [119] 2025.01 ResNet-101 Tr-Dec Introduced SECOND-CC, a large-scale dataset with semantic change masks and sentences describing changes. MModalCC integrates semantic masks through gated cross-attention mechanisms to facilitate captioning. [Link]
Mask Approx Net [121] 2025.02 ResNet Tr-Dec Proposed a diffusion-based method, Mask Approx Net, that transitions from feature learning to distribution learning. It comprises a multi-scale change detection module and a frequency-guided noise filter. [Link]
FST-Net [133] 2025.04 Segformer-B1 Tr-Dec FST-Net utilizes a Frequency-Spatial Fusion Module and a Spatial-Temporal Modeling module to adaptively separate structural signals from high-frequency noise and pseudo-changes.
@@ II-B2 Multi-task learning of Change Detection and Change Captioning
Change detection and change captioning address complementary aspects of remote sensing spatiotemporal vision-language understanding. Change detection focuses on producing pixel-level change masks to pinpoint changed areas, whereas change captioning aims to generate semantic-level descriptions that reveal object attributes and contextual relationships. Recognizing the intrinsic connection between these tasks, recent research has integrated them into a unified multi-task learning framework to enhance overall efficiency and accuracy [126, 128, 130, 131, 132]. Some representative methods are summarized in Table II.
One notable approach, Change-Agent [127], laid the groundwork by introducing a typical framework as shown in Fig. 4 (b). It employs a shared visual encoder to extract spatiotemporal features from remote sensing temporal images, followed by two task-specific branches: one for change detection and another for change captioning. In this framework, the change detection branch leverages multi-scale features to generate precise and detailed change masks, while the captioning branch typically utilizes deeper visual features to capture the semantics of the object changes. Such a design closely resembles single-task models in each branch yet benefits from joint optimization.
@@
change-aware VQA [136] 2022.07 CNN RNN First exploration of this task; constructed a multi-temporal image-question-answer triplet dataset and proposed a baseline method for this task.
CDVQA-Net [30] 2022.09 CNN RNN Built a large-scale public dataset and proposed an attention-based model encompassing multi-temporal feature encoding, fusion, multi-modal fusion, and an answer prediction head. [Link]
ChangeChat [137] 2024.09 CLIP-ViT Vicuna-v1.5 Constructed ChangeChat-87k dataset; Developed a LLM-based ChangeChat model to follow user’s complex instruction for multiple tasks, such as change classification and multi-turn conversation. [Link]
CDChat [138] 2024.09 CLIP ViT-L/14 Vicuna-v1.5 A contemporaneous work with ChangeChat, which adopts an architecture similar to LLava. It employs an MLP to bridge the CLIP image domain and the text domain of the LLM, and constructed a 19K instruction dataset. [Link]
TEOChat [94] 2024.10 CLIP ViT-L/14 LLaMA-2 Introduced TEOChat, a VLM tailored for multi-temporal earth observation. It demonstrated strong performance across multiple temporal tasks, outperforming the video-LLaVA model. [Link]
GeoLLaVA [139] 2024.10 Video encoder LLaVA-NeXT, Video-LLaVA Treats temporal images as video and utilizes a video encoder instead of the traditional image encoder; in addition, various efficient fine-tuning strategies for LLMs are investigated, such as LoRA and QLoRA. [Link]
VisTA [31] 2024.10 Shared CLIP Image Encoder CLIP Text Encoder Introduced Change Question Answering and Grounding task. Constructed the QAG-360K dataset with 360K triplets (questions, answers, and visual masks). Proposed a multi-stage semantic reasoning method. [Link]
RSUniVLM [140] 2024.12 Siglip-400m Qwen2-0.5B RSUniVLM is a unified, end-to-end VLM designed for multi-granularity understanding, following user’s instructions. It features a Granularity-oriented Mixture of Experts (Image-level, Region-level, and Pixel-level). [Link]
EarthDial [141] 2024.12 InternViT-300M Phi-3-mini EarthDial is a conversational VLM for multi-temporal, multi-spectral, and multi-resolution imagery. It introduces a large-scale instructional dataset with 11.11M pairs containing questions and corresponding responses. [Link]
UniRS [142] 2024.12 Siglip-400m Sheared-LLAMA-3B A unified VLM for understanding single images, dual-time images, and videos, which contains a change extraction module to handle temporal images, a prompt augmentation mechanism to leverage LLM knowledge. It achieves SOTA on RSCQA, captioning, and video scene classification. [Link]
DVLChat  [143] 2025.05 SAM Qwen2.5-VL Defined seven urban temporal tasks and evaluated 17 multimodal LLMs, revealing limitations in long-term temporal reasoning. Proposed a baseline model DVLChat capable of image-level QA and pixel-level segmentation, bridging dynamic urban insights through language-based interaction.
@@ II-B3 Remote Sensing Change Question Answering
In the visual encoding stage, the visual encoder is often employed to extract features from multi-temporal images, with the resulting features fused to capture dynamic visual information. In the question encoding stage, pre-trained language models (e.g., BERT [144] or GPT [145]) transform the user’s query into a semantic embedding suitable for model understanding. During multimodal interaction, attention mechanisms (e.g., self-attention and cross-attention) are widely applied to align visual features with the linguistic cues provided by the query, allowing the model to focus on the most relevant spatial regions. Finally, in the answer generation stage, the fused multimodal representation is converted into the language response.
@@ II-B4 Remote Sensing Text-driven Temporal Images Retrieval
One of the primary challenges in RSI-TIR is the issue of false negatives. Specifically, an image pair labeled as a negative sample in a training batch may actually be a positive sample that matches the query text, thus hindering effective model training. This problem is common in many contrastive learning tasks and can be addressed by some solutions [151, 152]. To mitigate the issue in RSI-TIR, Ferrod et al. [105] implemented two strategies: False Negative Elimination (FNE), which excludes potential false negatives from the loss computation to prevent interference, and False Negative Attraction (FNA), which re-labels these samples as positive to better align with the true relationships in the data. These techniques improve retrieval precision in complex spatiotemporal scenarios.
@@ II-B5 Remote Sensing Change Grounding
Methods such as TEOChat [94] and EarthDial [141] adopt bounding box outputs. For instance, TEOChat [94] use a model architecture inspired by LLaVA-1.5 [146], which utilizes a temporally shared ViT-L/14 to encode the temporal images. The resulting embeddings are mapped via an MLP and fed into LLaMA-2 [153], with the language model generating bounding box coordinates in textual format, thus effectively grounding the detected changes according to the input query.
On the other hand, methods that produce mask outputs include VisTA [31], ChangeChat [137], RSUniVLM [140], Falcon [154], and GeoRSMLLM [155]. For example, Li et al. introduced VisTA, a multi-task model designed for change question answering and grounding [31]. VisTA simultaneously generates a pixel-level change mask and a textual answer. The textual response is produced via a two-layer MLP, while the mask decoder employs two attention blocks, enabling the model to provide both semantic and visual explanations. This dual-output approach exemplifies a versatile solution for RS-CG tasks, enhancing the overall interpretability and precision of change localization in remote sensing imagery.
III Key Technologies in SpatioTemporal Visual-Language Models
III-A Preliminary Knowledge of Basic Models
III-A1 CNN
Convolutional Neural Networks (CNNs) are a class of deep learning models specifically designed to process grid-like data such as images. By leveraging local spatial correlations through convolutional operations, CNNs are capable of learning hierarchical feature representations in an end-to-end manner. A typical CNN architecture includes convolutional layers, non-linear activation functions (e.g., ReLU), pooling layers for spatial downsampling and normalization.
Due to their excellent capability in capturing local spatial patterns, CNNs have become fundamental in various remote sensing image interpretation tasks: 1) Scene Classification: To address object scale variation, Liu et al. [156] proposed a Scale-Robust CNN (SRSCNN) using random cropping to simulate scale differences and enhance feature robustness. 2) Object Detection: Li et al. [157] introduced LSKNet, which dynamically adjusts the receptive field via spatial selective kernels to better handle objects with varying spatial contexts. 3) Semantic Segmentation: Zhao et al. [158] developed SSAtNet with a ResNet-101 backbone and a pyramid attention pooling module for adaptive multiscale feature refinement. 4) Change Detection: Daudt et al. [65] pioneered the use of fully convolutional Siamese networks for change detection. In [159], Liu et al. provides a review of CNN in remote sensing.
III-A2 RNN and LSTM
Recurrent neural networks (RNNs) are designed to model sequential data by propagating historical information through recurrent hidden states, enabling the capture of temporal dependencies. At each time step, the hidden state is updated based on the current input $x_{t}$ and the previous hidden state $h_{t-1}$:
$h_{t}=\sigma(\mathbf{W_{h}}h_{t-1},\mathbf{W_{x}}x_{t})$
where $\sigma$ is the sigmoid activation function. However, standard RNNs often suffer from the vanishing gradient problem, limiting their ability to model long-term dependencies.
@@ III-A2 RNN and LSTM
To address this, Long Short-Term Memory (LSTM), a specialized type of RNNs, introduce memory cells and gating mechanisms to better control the flow of information, as shown in Fig. 5.
@@
where $x_{t}$ is the current input, $h_{t-1}$ is the previous hidden state.
@@
This architecture enables LSTM to dynamically capture long-term dependencies, making it widely applied in remote sensing, especially in sequence-related tasks: 1) Image Captioning: Qu et al. [161] first explored the “encoder-decoder” framework for caption generation and tried to use RNN and LSTM as language decoders. 2) Change Captioning: Chouaf and Hoxha et al. [95, 96] compared the performance of RNN and support vector machine (SVM) as language decoders for change captioning on two datasets. 3) Forecasting: Kim et al. [162] applied ConvLSTM for forecasting tropical cyclone paths.
III-A3 Transformer
@@
Figure 6: The architecture of Transformer [163].
Currently, the Transformer has been adapted in many fields, including NLP, computer vision [164, 165, 166, 167, 168], multimodality [169, 170, 171, 172, 173], medicine [174, 175], and remote sensing [176, 177, 178, 179]. In the past few years, the remote sensing community has also witnessed a significant growth in the use of transformer-based frameworks in many tasks: 1) Scene Classification: Ni et al. [180] proposed a second-order differentiable token transformer network with an efficient attention block and differentiable token compression to handle scene complexity and model redundancy. 2) Object Detection: Li et al. [181] introduced STDet, a sparse transformer-based detector with a local-to-global Transformer network (LGFormer) to extract object features, enhancing small object detection. 3) Semantic Segmentation: To capture detailed variations within small areas, Jing et al. [182] incorporated a hypergraph into the Transformer to learn general latent features as well as generate high-order visual representations by modeling correlations across multiscale features and local topology. 4) Change Detection: Yang et al. [183] proposed ConvFormer-CD, a CNN-Transformer hybrid model, to enhance global-local context awareness and mitigate pseudo-changes through Temporal Attention establishing cross-temporal semantic relationships between image pairs. Several surveys have reviewed Transformer applications in remote sensing, such as [176, 184].
III-A4 Mamba
@@
Mamba extends original SSMs by introducing a selective state space mechanism, allowing parameters like $\bar{\mathbf{B}}$, $\mathbf{C}$}, and $\mathbf{\Delta}$ to adapt dynamically to the input sequence, as shown in Fig. 7. This adaptability enables the model to selectively attend to different parts of the input, making it contextually aware of processing sequences in an efficient manner.
Figure 7: The architecture of Mamba [125].
Due to the global receptive field and linear computational complexity, Mamba [125] has demonstrated impressive performance across various tasks: 1) Scene Classification: Chen et al. [186] proposed a dynamic multi-path activation mechanism to augment Mamba’s capacity to model noncausal data and its sensitivity to spatial location. 2) Object Detection: Li et al. [157] proposed a mamba-based cross-modal fusion module to integrate RGB-Infrared features and improve discrimination between different classes. 3) Semantic Segmentation: Addressing semantic redundancy in multi-scale feature fusion, Wang et al. [187] proposed PyramidMamba, containing a plug-and-play decoder with dense spatial pyramid pooling and a pyramid fusion Mamba to reduce redundancy. 4) Change Detection: Chen et al. [87] introduced Mamba architecture into the change detection task and proposed three spatial-temporal relationship modeling mechanisms. More information about Mamba’s applications across various tasks can be found in [188, 189, 190, 191].
Figure 8: The key technologies can be applied across different remote sensing spatiotemporal vision-language tasks, including temporal image encoding, spatiotemporal visual feature fusion, language encoding, vision-language interaction, and language generation.